Overview of Status of Analysis, Design, Fabrication, Tests, etc.
Testing has been started and after talking to our sponsor, testing plans have been drawn out. Each test procedure is shown below and the number of times the test will be done is shown in parenthesis.
All tanks closed to find the COP of heat pump. There will be no outgoing flows from either cold or hot tank. (5)
Cold tank left open (cold water left running), and hot tank left at different constant temperatures. This measures the flow we need for cold tank to keep a certain degree in the hot tank constant. We will test for 43.3°C, 48.9°C, and 57.2°C. (110°F, 120°F, 135°F) (3+)
Constant cold temperature in cold tank and constant hot temperature in hot tank. Cold water will be calibrated to different temperatures to mimic different climate regions and will determine how much energy can be drawn from each calibration. We will attempt to test cold water temperatures at 10°C, 15.6°C (50°F, 60°F) (~6)
Electrical sensors have arrived and installed. But encountering problems working with them (not logging any data because of current wiring configuration of the system) and will be buying a Kil-A-Watt/extension cord/improvising either Wednesday or Thursday (2/22-23).
Improvements and additional sections added to the report and website (Sean, Cara)
Currently working on a new system model using immersion wells instead of our current temperature sensors for the residential models (Chris)
Labview simulation is working and assumptions seem reasonable when compared to the first test run. Assumptions are based on the specifications of the compressor. (Chris, Xavier)
Accomplishments from Previous Week
Electrical sensors have been picked up and talked to Steve Roberts on recommended installation methods. Decided to buy an extension cord and cut cord open rather than cut current cords connected to system. The transducer will wrap around extension cord's wire instead. (Xavier, Nadim, Cara)
Started testing procedure #1 and collecting data. (Everyone)
Talked to Don Lewis from Nyle and conducted further research on the calibration of the electrical hot water tank and the heat pump. Was previously unsure/encountered difficulties turning on the Nyle heat pump. (Sean, Xavier)
Added silicon sealant to sensor housing/plug and machined additional caps to block off possibility of plug bursting out of the tee pipes (Chris, Nadim)
Talked to Tom Chalfant about thermal wells for collecting temperature data for residential installations. Started a CAD model for the immersion wells. (Chris)
Collected missing receipts from sponsors and budget page has been updated. (Cara)
Goals for Next Week (list names after each item). Use specific and measurable objectives.
Try finishing all the runs for test #1 by end of this week and finish test #2 and #3 by end of next week - see above for test number references. (Everyone)
Astonish professor with our amazing report and website. (Everyone)
Finalize the immersion wells CAD model and make the part in machine shop (Chris, Nadim, Sean)
Figure out how to read flow meter data on HOBOlink's website by calling HOBO. Should be completed by end of this week (Cara)
Sponsor Comments from Last Meeting and Actions Taken to Address these Comments
Want to have at least one test where cold water is continuously run to figure out the energy available for the hot water. Also have cold water run at a 3:1 ratio with the hot water (i.e. 227.12 liters/60 gal cold to 75.7 liters/20 gal hot water). Added test procedure #3 in response to this comment
Instructor Comments from Last Meeting and Actions Taken to Address these Comments
Use the $240/year plan for HOBOlink data storage as backup plan. Proceed to use USB plug in for now to collect data.
Risks and Areas of Concern
Leakage on tee-pipes with temperature sensors, currently minor leak on sensor 4 & 2. However, 2 leaks infrequently. Will be remaking the housing for both 2 and 4 as soon as we buy more equipment.
Resources or Information Required but not Available
Schedule
Finish test #2 by Tuesday 2/28/12
Finish second draft of report by 3/1/12
Finish test #3 by Saturday 3/3/12
Budget (list amount spent and amount remaining)
Italics are items purchased in past. Items purchased this week are in normal print.
Progress on Report and Webpage
Final Design Section has been started and is being updated as final design changes/components as well as testing data comes up. (Sean)
Rough Draft sections edited based on instructor comments (Cara, Nadim)
Executive Summary in progress (Sean)
Budget section on website updated for 105.99 liter (20 gal) cold water tank + other piping equipment, and Nyle Heat pump. (Cara)